# Adaptive Mix for Semi-Supervised Medical Image Segmentation

### v1

@@ 4.2 Implementation Details
CPS (Chen et al., 2021) 86.23 76.22 11.68 3.65 88.72 80.01 7.49 1.91 CutMix Co-Training
@@
Data: In the 2D image segmentation tasks, all images are resized to $256\times 256$ for inference, while the outputs are recovered to the original size for evaluation. For 3D segmentation, we randomly crop $80\times 112\times 112\,(Depth\times Height\times Width)$ patches for training and iteratively crop patches using a sliding window strategy to obtain the final segmentation mask for evaluation.
4.3 Comparison with State of the Arts
We extensively compared our proposed AdaMix frameworks with state-of-the-art semi-supervised segmentation methods, which can be divided into four groups based on their perturbation strategies: 1) Multi-task perturbation (using signed distance maps (Li et al., 2020; Luo et al., 2021a) or multi-scale features (Luo et al., 2021b)): SASSNet (Li et al., 2020), URPC (Luo et al., 2021b), and DTC (Luo et al., 2021a); 2) Noise perturbation: UA-MT (Yu et al., 2019), MC-Net (Wu et al., 2021), MC-Net+ (Wu et al., 2022); 3) CutMix perturbation: FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), CPS (Chen et al., 2021); and 4) UMix and I-UMix perturbations: UCMT (Shen et al., 2023), and ABD (Chi et al., 2024).
Table 3: Comparison with state-of-the-art methods on the ISIC dataset with 5% $|\mathcal{D}|$ and 10% $|\mathcal{D}|$ labeled data. Note that all methods adopt U-Net as the backbone. The best and second-best results are highlighted in red and blue, respectively.
@@ 4.3 Comparison with State of the Arts
CPS (Chen et al., 2021) 83.03 74.06 14.67 3.80 84.99 76.79 11.13 2.20 CutMix Co-Training
@@ 4.3.1 Results on ACDC
Table 1 shows the results of our three AdaMix frameworks (i.e., AdaMix-ST, AdaMix-MT, and AdaMix-CT) and other semi-supervised segmentation methods on the ACDC dataset with 10% and 20% labeled data. All methods are evaluated on the average segmentation performance of left ventricular (LV) cavity, right ventricular (RV) cavity, and myocardium (MYO). The proposed AdaMix frameworks consistently outperform other approaches among all the metrics. Specifically, AdaMix-CT achieves the best performance of 89.19% and 89.83% in terms of DSC under 10% and 20% labeled data, respectively. Concerning the boundary-sensitive metrics 95HD and ASD, with the help of AdaMix, AdaMix-CT obtains the best values of 2.46 and 0.74 for the 10% labeled data. This implies that AdaMix can facilitate the model to delineate segmentation boundaries better. On the one hand, multi-task consistency-based approaches perform poorly in low data regions, i.e., 10% labeled data, indicating that the consistency regularization effects brought by auxiliary tasks, such as signed distance maps in SASSNet (Li et al., 2020) and DTC (Luo et al., 2021a), or deep supervision in URPC (Luo et al., 2021b), are limited. On the other hand, due to uncontrollable perturbation strength, the models equipped with either random noise or CutMix show a performance gap compared to our AdaMix framework. For instance, compared with the representative co-training-based model, MCNet+, our AdaMix-CT shows a performance gain of 1.48% and 1.32% in terms of DSC under 10% and 20% labeled data, respectively. Moreover, our AdaMix frameworks significantly outperform SSL methods with fixed mix-up rules, i.e., UCMT (Shen et al., 2023) and ABD (Chi et al., 2024). For example, compared to UCMT, AdaMix-CT obtains relative improvements of 2.62% in Dice similarity coefficient and 48.25% in average surface distance on the ACDC dataset with 10% labeled data. These results demonstrate the superiority of Adaptive Mix for consistency regularization in semi-supervised medical image segmentation.
@@ 4.3.2 Results on LA
We further evaluated the proposed method for left atrial segmentation in a 3D image segmentation scenario. Table 2 reports the results on the LA dataset with 10% (8 labeled samples) and 20% (16 labeled samples) labeled data. In general, our AdaMix frameworks substantially outperform all previous methods in terms of the four metrics. Specifically, the AdaMix frameworks show competitive performance in surface-related metrics, indicating that AdaMix, through adaptively adjusting perturbation degrees, generates reasonable perturbed images, thereby facilitating the frameworks’ advantages in 3D boundary segmentation. For example, AdaMix-CT obtains the best ASD values of 1.57 and 1.44 under 10% and 20% labeled data, respectively. These results prove the advantages of our method for semi-supervised volumetric medical image segmentation.
@@ 4.3.3 Results on ISIC
Unlike previous datasets for cardiac structure segmentation from MRI images, the ISIC dataset involves skin lesion segmentation from dermoscopy images, which includes heterogeneous segmentation objects. The results are reported in Table 3. It can be observed that in comparison with the methods utilizing CutMix (Yun et al., 2019) perturbation (e.g., FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021)) or UMix (Shen et al., 2023) perturbation (e.g., UCMT (Shen et al., 2023) and ABD (Chen et al., 2021)), our AdaMix frameworks demonstrate compelling performance. For instance, compared with UCMT (Shen et al., 2023) and ABD (Chi et al., 2024), AdaMix-CT shows performance gains of 1.64% and 1.26% in terms of DSC respectively under 5% labeled data. Remarkably, in light of boundary-sensitive metrics, i.e., 95HD and ASD, the proposed methods obtain the best scores due to effective training on the challenging target locations and boundaries in the perturbed images provided by AdaMix. These results further demonstrate the superiority of our method for heterogeneous lesion segmentation.
Table 4: Ablation study of the proposed AdaMix-ST framework on the ACDC val set under 10% labeled data. Sup: the supervised baseline. Note that without the self-paced mask, AdaMix is performed according to a fixed rule, i.e., replacing low-confidence patches with high-confidence ones. The best results are highlighted in bold.
@@
Figure 4: Investigation of the effect of patch size $S$ and No. patches $K$ on AdaMix using the ACDC val set with 10% labeled data.
Remark One can observe from Table 1, Table 2, and Table 3 that the results obtained by the three AdaMix frameworks are comparable, validating the flexibility of AdaMix to act as an effective plug-in-play module in SSL paradigms. In contrast, the methods with different perturbation strategies show obvious performance discrepancies. For example, MT and AdaMix-MT exhibit a performance gap of 4.05% in terms of DSC on the LA dataset with 10% labeled data, while the margin between MT and CPS is only 0.08%. The results suggest that perturbation strategies are more essential than learning paradigms for enhancing the models’ performance in consistency regularization-based methods. This phenomenon can be attributed to the data-driven nature: DNNs are data-driven, making the quality of training data more crucial for accurately fitting the real data distribution in the underlying task, where the quality of training data can be considered as the degree to which the training data distribution represents the real data distribution.
@@ 4.4 Ablation Study
In the following, we conducted extensive ablation experiments to investigate the contribution of each component of the proposed method in detail [Table 4]. Specifically, we took our AdaMix-ST as an example and examined the effects of AdaMix by considering three versions: 1) AdaMix-ST (V1): AdaMix-ST without the self-paced weight, 2) AdaMix-ST (V2): AdaMix-ST without the self-paced mask, and 3) AdaMix-ST (V3): the final method. Furthermore, we explored in-depth the influence of hyperparameters, i.e., the maximum number of mix-up patches $K$ and the patch size $S$ [Fig. 4].
@@ 4.4.1 Improvement over the Supervised Baseline
Compared to the supervised baseline, our final method (i.e., AdaMix-ST (V3)) achieves considerable relative improvements of 5.97%, 9.99%, 31.80%, and 58.25%, in terms of DSC, Jaccard, 95HD, and ASD, respectively. These results demonstrate the effectiveness of our AdaMix frameworks for semi-supervised medical image segmentation.
@@ 4.4.2 Effectiveness of Each Component
It can be observed that when a self-paced mask is introduced into the model to determine the mix-up rule based on the model’s learning state, the results of AdaMix-ST (V1) are higher than those of SSL-ST, for example, with 2.08% and 1.06 improvement in terms of DSC and ASD, respectively.
@@
We can conclude that our self-paced curriculum for determining both the mix-up rule and the number of patches is beneficial for semi-supervised segmentation. Building upon the self-paced curriculum, our final model, AdaMix-ST (V3), shows significant improvements over the semi-supervised baseline in global segmentation metrics (DSC and Jaccard) and boundary-sensitive metrics (95HD and ASD), achieving the best performance with 88.86% DSC, 81.03% Jaccard, 4.46 95HD, and 1.24 ASD.
Figure 5: Visualization of the AdaMix procedure during training: AdaMix initially generates perturbed samples containing more high-confidence regions, and then gradually incorporates low-confidence regions into the perturbed images as training progresses (green regions: the high-confidence patches from auxiliary images; red regions: the low-confidence patches from auxiliary images). According to the confidence maps, high-confidence regions generally appear in the background, while low-confidence ones tend to appear at the boundaries of the segmentation targets.
4.5 Analysis of Hyperparameters
@@
We investigated the influence of the mix-up patch size $S$ and the maximum number of mix-up patches $K$ on AdaMix, using the ACDC dataset with 10% labeled data. The results are shown in Fig. 4, where each subplot illustrates how the DSC changes with varying patch sizes when the maximum number of mix-up patches is fixed. In general, AdaMix is not sensitive to changes in $K$, whereas variations in $S$ cause slight performance fluctuations. Specifically, increasing the patch size marginally improves performance, with the highest DSC score achieved when $S=32$. The main reasons behind this phenomenon are: 1) the proposed self-paced learning strategy can adaptively adjust the number of mix-up patches based on the model’s learning state, where $K$ only determines the maximum number of mix-up patches, and 2) patches that are too small or too large inevitably introduce noise to the images. Small patches fail to provide complete anatomical information, while large patches disrupt the anatomical structure of original images. Moreover, it’s surprising that the impact of the mix-up patch size K on the variation of segmentation performance is more trivial when the maximum number of mix-up patches $K=32$. We speculate that with a large number of mix-up patches, AdaMix can control a wider range of mix-up intensities. Consequently, it exhibits less sensitivity to hyperparameters, making the algorithm more robust.
4.6 Analysis of Adaptive Mix Samples

### v2

@@ 4.3 Comparison with State of the Arts
CPS (Chen et al., 2021b) 86.23 76.22 11.68 3.65 88.72 80.01 7.49 1.91 CutMix Co-Training
@@
Table 3: Comparison with state-of-the-art methods on the ISIC dataset with 5% $|\mathcal{D}|$ and 10% $|\mathcal{D}|$ labeled data. Note that all methods adopt U-Net as the backbone. The best and second-best results are highlighted in red and blue, respectively. indicates statistically significant improvements ($p$-value $<0.05$ based on a paired t-test) of our AdaMix frameworks over ABD, MT, and UCMT, respectively.
@@
CPS (Chen et al., 2021b) 83.03 74.06 14.67 3.80 84.99 76.79 11.13 2.20 CutMix Co-Training
@@ 4.3.1 Results on ACDC
Table 1 shows the results of our three AdaMix frameworks (i.e., AdaMix-ST, AdaMix-MT, and AdaMix-CT) and the compared methods on the ACDC dataset with 10% and 20% labeled data. All the methods are evaluated on the average segmentation performance of left ventricular (LV) cavity, right ventricular (RV) cavity, and myocardium (MYO). The proposed AdaMix frameworks consistently outperform other approaches among all the metrics. Specifically, AdaMix-CT achieves the best performance of 89.19% and 89.83% in terms of DSC under 10% and 20% labeled data, respectively. Concerning the boundary-sensitive metrics 95HD and ASD, with the help of AdaMix, AdaMix-CT obtains the best values of 2.46 and 0.74 for the 10% labeled data. This implies that AdaMix can facilitate the model to delineate segmentation boundaries better. On the one hand, multi-task consistency-based approaches perform poorly in low data regions, i.e., 10% labeled data, indicating that the consistency regularization effects brought by auxiliary tasks, such as signed distance maps in SASSNet (Li et al., 2020) and DTC (Luo et al., 2021a), or deep supervision in URPC (Luo et al., 2021b), are limited. On the other hand, due to uncontrollable perturbation strength, the models equipped with either random noise or CutMix show a performance gap compared to our AdaMix framework. For instance, compared with the representative co-training-based model, MCNet+, our AdaMix-CT shows a performance gain of 2.09% and 1.32% in terms of DSC under 10% and 20% labeled data, respectively. Moreover, our AdaMix frameworks significantly outperform the SSL methods with fixed mix-up rules, i.e., UCMT (Shen et al., 2023) and ABD (Chi et al., 2024). For example, compared with UCMT, AdaMix-CT obtains relative improvements of 2.62% in Dice similarity coefficient and 48.25% in average surface distance on the ACDC dataset with 10% labeled data. These results demonstrate the superiority of Adaptive Mix for consistency regularization in semi-supervised medical image segmentation.
@@ 4.3.2 Results on LA
We further evaluated the proposed method on a 3D image segmentation scenario for left atrial segmentation. Table 2 reports the results on the LA dataset with 10% (8 labeled samples) and 20% (16 labeled samples) labeled data. In general, our AdaMix frameworks substantially outperform all the compared methods in terms of the four metrics. Specifically, the AdaMix frameworks show competitive performance in surface-related metrics, indicating that AdaMix, through adaptively adjusting perturbation degrees, generates appropriate perturbed images, thereby facilitating the frameworks’ advantages in 3D boundary segmentation. For example, AdaMix-CT obtains the best ASD values of 1.57 and 1.44 under 10% and 20% labeled data, respectively. These results prove the advantages of our method for semi-supervised volumetric medical image segmentation.
@@ 4.3.3 Results on ISIC
Unlike previous datasets for cardiac structure segmentation from MRI images, the ISIC dataset involves skin lesion segmentation from dermoscopy images, which includes heterogeneous segmentation objects. The results are reported in Table 3. It can be observed that in comparison with the methods utilizing CutMix (Yun et al., 2019) perturbation (e.g., FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021b)) or UMix (Shen et al., 2023) perturbation (e.g., UCMT (Shen et al., 2023) and ABD (Chen et al., 2021b)), our AdaMix frameworks demonstrate compelling performance. For instance, compared with UCMT (Shen et al., 2023) and ABD (Chi et al., 2024), AdaMix-CT shows performance gains of 1.64% and 1.26% in terms of DSC respectively under 5% labeled data. Remarkably, in light of boundary-sensitive metrics, i.e., 95HD and ASD, the proposed methods obtain the best scores due to sufficient training on the challenging target locations and boundaries in the perturbed images provided by AdaMix. These results further demonstrate the superiority of our method for heterogeneous lesion segmentation.
Remark. One can observe from Table 1, Table 2, and Table 3 that the results obtained by the three AdaMix frameworks are comparable, validating the flexibility of AdaMix to act as an effective plug-in-play module in SSL paradigms. In contrast, the methods with different perturbation strategies show significant performance discrepancies. For example, MT and AdaMix-MT exhibit a performance gap of 4.05% in terms of DSC on the LA dataset with 10% labeled data, while the margin between MT and CPS is only 0.08%. The results suggest that perturbation strategies are more critical than learning paradigms for enhancing the models’ performance in consistency regularization-based methods. This phenomenon can be attributed to the data-driven nature: DNNs are data-driven, making the quality of training data more crucial for accurately fitting the real data distribution in the underlying task, where the quality of training data can be considered as the degree to which the training data distribution represents the real data distribution.
4.3.4 Qualitative Results
As shown in Fig. 4, we visualize some segmentation examples on the ACDC, LA, and ISIC datasets. Specifically, the supervised baseline is prone to under-segment some lesion regions due to the limited number of labeled data. FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021b) (all using CutMix (Yun et al., 2019) as the strong perturbation) incorrectly classify many boundary pixels, mainly due to noise introduced by random perturbations during training. Similarly, UCMT (Shen et al., 2023) and ABD (Chi et al., 2024), which use UMix (Shen et al., 2023) as the strong perturbation, also exhibit inferior performance on segmentation boundaries. These results can be attributed to the limited effectiveness of consistency regularization. In contrast, our AdaMix-ST exhibits smoother and clearer boundaries compared to other methods. The finer segmentation results can be ascribed to the AdaMix algorithm, which adaptively generates perturbed images in a self-paced learning manner.
Table 4: Ablation study of the AdaMix-ST framework on the ACDC val set under 10% labeled data. Sup: the supervised baseline. Note that without the self-paced mask, AdaMix is performed according to a fixed rule, i.e., replacing low-confidence patches with high-confidence ones. The best results are highlighted in bold.
@@ 4.3.4 Qualitative Results
Figure 5: Investigation of the effect of patch size $S$ and maximum No. patches $K$ for AdaMix on the ACDC val set with 10% labeled data.
Figure 6: Analysis of the effect of age parameter functions for AdaMix on the ACDC dataset with 10% labeled data. Left: the age parameter function curves; middle: the training loss curves of models with different age parameter functions; right: the corresponding segmentation performance on the ACDC val set.
@@ 4.4 Ablation Study
In the following, we conducted extensive ablation experiments to investigate the contribution of each component of the proposed method in detail [Table 4]. Specifically, we took our AdaMix-ST as an example and examined the effects of AdaMix by considering three versions: 1) AdaMix-ST (V1): AdaMix-ST without the self-paced weight, 2) AdaMix-ST (V2): AdaMix-ST without the self-paced mask, and 3) AdaMix-ST (V3): the final method. Meanwhile, we explored in-depth the influence of hyperparameters, i.e., the maximum number of mix-up patches $K$ and the patch size $S$ [Fig. 5], as well as the age parameter $\lambda$ [Fig. 6]. Moreover, we also provide a theoretical and experimental analysis of the effect of AdaMix on training convergence [Fig. 8], and visualize several AdaMix examples during training [Fig. 9].
@@ 4.4.1 Improvement over the Supervised Baseline
Compared with the supervised baseline, our final method (i.e., AdaMix-ST (V3)) achieves considerable relative improvements of 5.97%, 9.99%, 31.80%, and 58.25%, in terms of DSC, Jaccard, 95HD, and ASD, respectively. These results demonstrate the effectiveness of our AdaMix frameworks for semi-supervised medical image segmentation.
@@ 4.4.2 Effectiveness of Each Component
It can be observed that when a self-paced mask is introduced into the model to determine the mix-up rule based on the model’s learning state, the results of AdaMix-ST (V1) are higher than those of SSL-ST, e.g., with 2.08% and 1.06 improvement in DSC and ASD, respectively.
@@
We can conclude that the self-paced curriculum for determining the mix-up rule and the number of patches is beneficial for semi-supervised segmentation. Building upon the self-paced curriculum, our final model, AdaMix-ST (V3), shows significant improvements over the semi-supervised baseline in both global segmentation metrics (DSC and Jaccard) and boundary-sensitive metrics (95HD and ASD), achieving the best performance with 88.86% DSC, 81.03% Jaccard, 4.46 95HD, and 1.24 ASD.
Figure 7: Gradient variances of models trained with and without AdaMix on the ACDC, LA, and ISIC datasets. Each point on the curves represents the norm of the batch-wise gradient variances for the entire training set.
Figure 8: Training and validation losses of CutMix, UMix, I-UMix, and the proposed AdaMix on the (a) ACDC, (b) LA, and (c) ISIC datasets, respectively.
4.4.3 Analysis of Hyperparameters
@@
We investigated the influence of the mix-up patch size $S$ and the maximum number of mix-up patches $K$ on AdaMix, using the ACDC val set with 10% labeled data. The results are shown in Fig. 5, where each subplot illustrates how the DSC changes with varying patch sizes when the maximum number of mix-up patches is fixed. In general, AdaMix is not sensitive to changes in $K$, whereas variations in $S$ cause slight performance fluctuations. Specifically, increasing the patch size marginally improves performance, with the highest DSC score achieved when $S=32$. The main reasons behind this phenomenon are: 1) the proposed self-paced learning strategy can adaptively adjust the number of mix-up patches based on the model’s learning state, where $K$ only determines the maximum number of mix-up patches, and 2) patches that are too small or too large inevitably introduce noise to the images; small patches fail to provide complete anatomical information, while large patches disrupt the anatomical structure of original images. Moreover, it is surprising that the impact of the mix-up patch size $S$ on the variation in performance is more trivial when the maximum number of mix-up patches $K=32$. We believe that with a large number of mix-up patches, AdaMix can control a wider range of mix-up degrees. Consequently, it exhibits less sensitivity to hyperparameters, making the algorithm more robust.
4.4.4 Analysis of Age Parameter Functions
We explored the impact of different age parameter functions for AdaMix on the ACDC dataset with 10% labeled data. As shown in Fig. 6, while different function shapes lead to variations in the adaptive perturbation procedure, all models with these functions converge to minima during training. However, these variations in adaptive perturbation impact the segmentation performance of the trained models. Notably, the age parameter $\lambda$ increasing with a Gaussian ramp-up function contributes to the highest DSC score, compared with the other functions. These results suggest that the Gaussian function curve aligns better with the model’s state in the training process and validates our hypothesis that the model’s performance progressively improves during training.
Figure 9: Visualization of the AdaMix procedure during training: AdaMix initially generates perturbed samples containing more high-confidence regions, and then gradually incorporates low-confidence regions into the perturbed images as training progresses (green regions: the high-confidence patches from auxiliary images; red regions: the low-confidence patches from auxiliary images). According to the confidence maps, high-confidence regions generally appear in the background, while low-confidence ones tend to occur at the boundaries of the segmentation targets.
4.4.5 Convergence Analysis
Based on the optimization theory for large-scale machine learning (Bottou et al., 2018), we first provide a theoretical analysis that the AdaMix algorithm can generate perturbed samples to reduce gradient variance for facilitating faster convergence of the segmentation model training.
Let $J(\theta)$ be the empirical risk over the dataset $\mathcal{D}$ for the segmentation loss function $\mathcal{L}(\theta)$ with the model $f(\cdot;\theta)$. We assume that the loss function is $\ell$-gradient Lipschitz and that the gradient estimate has bounded variance. Based on the $\ell$-smooth ($\ell$-gradient Lipschitz) and the bounded gradient variance assumptions, we have the following convergence theorem.
Theorem 1
Convergence theorem. Using the stochastic gradient descent with a learning rate $\eta_{t}=\frac{\eta_{0}}{t}$ at the $t^{th}$ iteration ($\eta_{0}$ refers to the initial learning rate) and the total optimization steps $T$, the convergence rate satisfies:
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\|J(\theta_{t})\|^{2}]\leq\frac{2(J(\theta_{0})-J^{*}))}{\eta_{0}\sqrt{T}}+\frac{\eta_{0}\ell\sigma^{2}}{\sqrt{T}}$
where the term $\frac{\eta_{0}\ell\sigma^{2}}{\sqrt{T}}$ dominates the bound for large $T$.
This theorem implies that a smaller $\sigma^{2}$ ensures faster convergence to a stationary point. Since AdaMix performs adaptive patch-wise mix-up between two images and its loss function is the weighted average over these regions, the resulting loss function for AdaMix samples still satisfies $\ell$-smooth, and its Lipschitz constant remains $\ell$. Mix-up operations between different images generate the synthetic samples, which lie in the neighborhood of the original images, leading to a smoother estimate of the gradient, i.e. a lower gradient variance $\sigma^{2}_{x}\leq\sigma^{2}$  (Zhang et al., 2017; Yun et al., 2019).
We conducted experimental validation to support this argument. To this end, we compared the gradient variances of models trained with and without AdaMix on the ACDC, LA, and ISIC datasets. As shown in Fig. 7, the model trained with AdaMix has lower gradient variances, indicating a more stable training process. Notably, AdaMix maintains consistently smaller gradient variances throughout the training process, except for some fluctuations observed during the mid-training phase on the LA dataset. This result corroborates our theoretical analysis, confirming that the AdaMix algorithm reduces gradient variance.
We visualize the training and validation loss curves of CutMix, UMix, I-UMix, and our AdaMix on the ACDC, LA, and ISIC datasets in Fig. 8, respectively. According to the training loss curves, these methods generally converge to their minima. It can be observed that both CutMix and I-UMix exhibit more dramatic oscillations during the training process, especially on the ISIC dataset. This is because the random mixing operation or the mixing with low-confidence regions tends to generate perturbed samples that deviate from the model’s current state. Meanwhile, UMix leads to a faster decrease in training loss but with a relatively inferior generalization performance on the validation sets, as mixing with high-confidence regions produces easier samples. In contrast, AdaMix, with its adaptive perturbation strength, ensures a more stable training process. On the other hand, the lower validation loss of AdaMix indicates that it leads to a more robust trained model, i.e., against overfitting and with better generalization capability. In summary, through the empirical experiments, we demonstrate that the AdaMix algorithm reduces gradient variance, leading to faster training convergence.
4.4.6 Analysis of Adaptive Mix Samples